THE STUDY OF NOBLE GAS - NOBLE METAL HALIDE INTERACTIONS: FOURIER TRANSFORM MICROWAVE SPECTROSCOPY OF XeCuF

\underline{JULIE M. MICHAUD} AND MICHAEL C. L. GERRY, Department of Chemistry, The University of British Columbia, 2036 Main Mall, Vancouver, B.C., Canada V6T 1Z1.

High resolution Fourier transform microwave (FTMW) spectroscopy has been used to measure the rotational spectra of 7 isotopomers of the complex XeCuF. The molecules were prepared by laser ablation of Cu metal in the presence of xenon and SF_{6}. The resulting molecules were stabilized in a supersonic jet of Ar injected into the cavity of the FTMW spectrometer. The spectroscopic constants obtained include the rotational constants and centrifugal distortion constants (B_{0}, D_{0}). Nuclear quadrupole coupling constants obtained for Cu and ^{131}Xe indicate significant reorganization of the electron distribution on complex formation. The Xe-Cu distance obtained from the experimental data is small compared to the sum of the Xe van der Waals radius and the Cu^{+} ionic radius. The nature of the Xe-CuF interaction will be discussed.